# Provable Mixed-Noise Learning with Flow-Matching

### v1

@@ Abstract
We study Bayesian inverse problems with mixed noise, modeled as a combination of additive and multiplicative Gaussian components. While traditional inference methods often assume fixed or known noise characteristics, real-world applications, particularly in physics and chemistry, frequently involve noise with unknown and heterogeneous structure. Motivated by recent advances in flow-based generative modeling, we propose a novel inference framework based on conditional flow matching embedded within an Expectation-Maximization (EM) algorithm to jointly estimate posterior samplers and noise parameters. To enable high-dimensional inference and improve scalability, we use simulation-free ODE-based flow matching as the generative model in the E-step of the EM algorithm. We prove that, under suitable assumptions, the EM updates converge to the true noise parameters in the population limit of infinite observations. Our numerical results illustrate the effectiveness of combining EM inference with flow matching for mixed-noise Bayesian inverse problems.
@@ 1 Introduction
where $F\colon\mathbb{R}^{m}\to\mathbb{R}^{n}$ is a known possibly nonlinear forward operator and $\eta_{\theta}$ represents random noise depending on parameters $\theta$. In many real-world settings, the noise is not known, but depends on auxiliary noise parameters $\theta$, which are in our case to scalars $(a,b)$. We focus on so-called mixed noise, which is a sum of additive and multiplicative Gaussian noise, i.e.,
@@
As many physical phenomena (or approximations thereof) do not usually exhibit Gaussian errors, this often yields a more flexible modelling approach. It is present in optics [14, 13] and also biology [22].
In [11], an iterative procedure was introduced to jointly update normalizing flows and noise parameters, formulated within the Expectation-Maximization (EM) framework [8]. This approach can be interpreted as a conditional extension of the DeepGEM algorithm [9]. However, the only convergence guarantee provided was the non-decreasing behavior of the marginal log-likelihood across iterations. Moreover, the method relies on discrete normalizing flows [23], which have recently been surpassed in performance and scalability by ODE-based normalizing flows [6], particularly those trained in a simulation-free manner via flow matching [18, 20]. Flow Matching has shown great success, in particular in image generation and regularization [21], and is a variant of score-based diffusion [24] which does not rely on a standard Gaussian latent. This enables interesting applications [26]. Further, we can also make the latent distribution dependent on the data distribution, which yields optimal transport inspired flow matching approaches [5, 29].
@@
  • formulate an EM algorithm where the E-step is realized via conditional flow matching, a modern generative modeling technique based on score-based ODEs;
@@
  • establish consistency guarantees: in the population limit of infinitely many measurements generated under a true noise parameter $\theta^{*}$, the EM procedure converges to $\theta^{*}$ under suitable regularity assumptions.
@@
We denote by $X\in\mathbb{R}^{m}$ the latent random variable, by $Y\in\mathbb{R}^{n}$ the observation and $\Theta\subset\mathbb{R}^{d}$ a nonempty, compact and convex parameter space. We denote by $\|\cdot\|$ the Euclidean norm on $\mathbb{R}^{d}$. For a function $\mathcal{Q}\colon\Theta\times\Theta\to\mathbb{R}$ we denote by $\nabla_{1}$, the gradient with respect to the first input of a function with two arguments, e.g. $\nabla_{1}\mathcal{Q}$.
Assuming existence (cp. [27]), we denote by $p_{Y_{\theta}}$ the density of the marginal likelihood of $Y$ given $\theta$, by $p_{X|Y_{\theta}=y}$ the corresponding posterior density and by $p_{X,Y_{\theta}}$ the joint density. The corresponding distributions are distinguished using capitalization, e.g. $P_{X|Y_{\theta}=y}$ for the posterior distribution. We denote by $\mathrm{pdf}(\mathbb{R}^{d})$ the set of all Lebesgue densities on $\mathbb{R}^{d}$, i.e.,
@@
whenever $\mu$ is absolutely continuous with respect to $\nu$. For measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$ with Lebesque densities $f,g\in\mathrm{pdf}(\mathbb{R}^{d})$ we may use abuse of notation and define
@@
Outline.
The manuscript on hand is organized as follows. We start giving the relevant background information on the general EM algorithm and Flow Matching in Section˜2.1 and Section˜2.2 respectively. Next, we recall the convergence analysis of the so-called population EM operator in Section˜3 and extend the theory in the follow up subsections Section˜3.3 and Section˜3.4. Then, we will verify the new framework for a class of mixed noise models in Section˜4. Finally, we illustrate the proposed method of mixed-noise learning with flow-matching in the numerical section for MNIST denoising in Section˜5.1 and for a reconstruction task with underlying partial differential equation (PDE) in Section˜5.2.
@@ 2.1 EM Algorithm
For the E-step, classical approaches include MCMC-based posterior sampling. However, such methods are computationally demanding, especially when repeated for each observation. Alternatively, we adopt a conditional generative modeling perspective, where a single shared model is trained to approximate the posterior distribution for all observations. This amortized setup enables efficient inference across the dataset and allows parameter sharing across different $y_{i}$. While [11] relied on conditional normalizing flows in the spirit of [9], we employ conditional flow matching as introduced in [1, 18, 19], which provides improved stability and scalability in high-dimensional settings.
@@ 2.2 Flow Matching
We now detail the concrete realization of the E-step in our EM algorithm using conditional flow matching.
E-step via Conditional Flow Matching.
The E-step aims to approximate the posterior distributions
@@
for each observation $y_{i}$, given current parameter estimates $\theta^{(k)}$. In [11], these posteriors were approximated using conditional normalizing flows, extending the DeepGEM framework [9] to the case of unknown noise parameters.
@@ 3 Convergence Analysis of the Population EM Operator
We start by recalling the convergence results for the population EM algorithm from [2]. For $\theta\in\Theta$ consider a parametric family of distributions $P_{Y_{\theta}}$ with density $p_{Y_{\theta}}$. We assume that the observed data are generated according to a distribution $P_{\theta^{*}}=P_{Y_{\theta^{*}}}$ with density $p_{\theta^{*}}=p_{Y_{\theta^{*}}}$ corresponding to a true but unknown parameter $\theta^{*}\in\Theta$. More precisely, we assume that $\theta^{*}$ is the unique maximizer of the population log-likelihood
$\theta^{*}=\operatorname*{arg\,max}_{\theta\in\Theta}\,\mathbb{E}_{y\sim p_{\theta^{*}}}[\log p_{Y_{\theta}}(y)].$
@@
$\mathcal{Q}(\theta,\hat{\theta}):=\mathbb{E}_{y\sim p_{\theta^{*}}}\,\mathbb{E}_{x\sim p_{X\mid Y_{\hat{\theta}}=y}}\left[\log p_{X,Y_{\theta}}(x,y)\right].$
@@ 3.1 Sufficient Assumptions
In this subsection, we are interested to derive sufficient conditions that yield locally strong concavity and first order stability on one hand for the case of population EM iteration to imply Theorem˜1 and locally strong concavity, $\gamma$-gradient smoothness and $\mu$-smoothness for the case of gradient-ascent based iteration to imply Theorem˜2.
@@
    (Smoothness and Integrability of Log-Likelihood) The map $\theta\mapsto\log p_{X,Y_{\theta}}$is two times continuously differentiable in $B_{\varepsilon}(\theta^{*})$ for all $(x,y)\in\mathbb{R}^{m\times n}$ and some $\varepsilon>0$ and there exist functions $G,H:\mathbb{R}^{m\times n}\to\mathbb{R}$ such that for all $(x,y)\in\mathbb{R}^{m\times n}$ and all $\theta\in B_{\varepsilon}(\theta^{*})$ it holds
@@
    $\int_{\mathbb{R}^{m\times n}}G(x,y)\,p_{X|Y_{\theta}=y}\,p_{\theta^{*}}\text{ d}x\text{ d}y<C,\quad\int_{\mathbb{R}^{m\times n}}H(x,y)\,p_{X|Y_{\theta}=y}\,p_{\theta^{*}}\text{ d}x\text{ d}y<C.$
@@
    (Local identifiability of $\theta^{\ast}$) The Fisher information
    $\mathcal{I}(\theta^{\ast})\coloneqq-\mathbb{E}_{x,y}[\nabla_{1}´{\theta}^{2}\log p_{X,Y_{\theta^{\ast}}}]$
    satisfies $\mathcal{I}(\theta^{*})\geq\lambda_{\ast}I_{d}$, for some $\lambda_{\ast}>0$ and further also satisfies $\mathcal{I}(\theta^{*})\leq\lambda_{2}I_{d}$ for a $\lambda_{2}>0$.
@@
    The map $\theta\mapsto p_{X|Y_{\theta}=y}$ is locally $L(y)$- Lipschitz in the Wasserstein-1 distance, i.e. for all $(x,y)\in\mathbb{R}^{m\times n}$
@@
    $\left\|\nabla_{\theta}\log p_{X,Y_{\theta}}(x_{1},y)-\nabla_{\theta}\log p_{X,Y_{\theta}}(x_{2},y)\right\|\leq L_{g}(y,\theta)\|x_{1}-x_{2}\|\quad\forall y\in\mathbb{R}^{n}.$
@@
    $\gamma=\gamma(\varepsilon)=\sup\limits_{\theta\in B_{\varepsilon}(\theta^{*})}\mathbb{E}_{y\sim p_{\theta^{*}}}[L(y)L_{g}(\theta,y)]<\infty.$
@@
    $\gamma=\gamma(\varepsilon)=\sup\limits_{\theta\in B_{\varepsilon}(\theta^{*})}\mathbb{E}_{y\sim p_{\theta^{*}}}[L(y)L_{g}({M}(\theta),y)]<\infty.$
@@ 3.2 FOS Framework
Let (A1) and (A2) be fulfilled. Then there exist $\varepsilon>0$ and $\lambda=\lambda(\varepsilon,\lambda^{\ast})>0$ such that $\mathcal{Q}(\cdot,\theta^{\ast})$ is strong $\lambda$-concave on $B_{\varepsilon}(\theta^{*})$.
@@
Under regularity assumptions of (A1) we can switch the following differentiation and integration, and applying subsequently (A2), we get
$\nabla_{1}^{2}\mathcal{Q}(\theta^{\ast},\theta^{\ast})=\mathbb{E}_{y\sim p_{\theta^{*}}}\mathbb{E}_{x\sim P_{X|Y_{\theta^{*}}=y}}\left[\nabla_{\theta}^{2}\log p_{X,Y_{\theta^{\ast}}}(x,y)\right]=-\mathcal{I}(\theta^{\ast})\leq-\lambda_{\ast}I_{d}.\\$ (5)
@@
for some $\lambda=\lambda(\varepsilon,\lambda^{\ast})>0$, which yields the strong $\lambda$-concavity $\mathcal{Q}(\cdot,\theta^{\ast})$. ∎
@@
$\nabla_{1}\mathcal{Q}(\theta,\hat{\theta})-\nabla_{1}\mathcal{Q}(\theta,\theta^{*})$ $=\mathbb{E}_{y\sim p_{\theta^{*}}}\Big{[}\mathbb{E}_{x\sim P_{X|Y_{\hat{\theta}}=y}}\left[\nabla_{\theta}\log p_{X,Y_{\theta}}(x,y)\right]$
@@
$=\mathbb{E}_{y\sim p_{\theta^{*}}}\Big{[}\mathbb{E}_{x\sim P_{X|Y_{\hat{\theta}}=y}}\left[\phi_{y,\theta}(x)\right]-\mathbb{E}_{x\sim P_{X|Y_{\theta^{*}}=y}}\left[\phi_{y,\theta}(x)\right]\Big{]}.$
@@
Integrating over $y\sim p_{\theta^{*}}$, we obtain
$\left\|\nabla_{1}\mathcal{Q}(\theta,\hat{\theta})-\nabla_{1}\mathcal{Q}(\theta,\theta^{*})\right\|\leq\mathbb{E}_{y\sim p_{\theta^{*}}}[L(y)L_{g}(\theta,y)]\|\hat{\theta}-\theta^{\ast}\|.$
@@
$\left\|\nabla_{1}\mathcal{Q}(M(\hat{\theta}),\hat{\theta})-\nabla_{1}\mathcal{Q}(M(\hat{\theta}),\theta^{*})\right\|$ $\leq\mathbb{E}_{y\sim p_{\theta^{*}}}[L(y)L_{g}(M(\hat{\theta}),y)]\|\hat{\theta}-\theta^{\ast}\|$
@@
Hence, we get slightly different assumptions: The score $\nabla_{\theta}\log p_{X,Y_{\theta}}$ now only needs to be $L^{\infty}$, whereas the Lipschitz continuity bound is a bit harder to satisfy.
Note that in order to apply the convergence theorem, we must assume that $0\leq\gamma<\lambda$. However, there are two main reasons why we prefer the following gradient-based framework over the classical EM convergence theory relying on the FOS condition. First, the FOS condition involves the population EM operator $M$, whose analytical form is often intractable and difficult to estimate numerically, especially in complex models. Second, Lemma 8 provides a slightly improved convergence guarantee: instead of requiring $\gamma<\lambda$, it suffices to assume $\gamma<c$, where $c$ is the strong convexity constant of the Q-function. This is a strictly weaker condition in non-degenerate cases. Therefore, we adopt the gradient-based framework, which is both more natural for our approximate M-step and yields more accessible convergence criteria.

### v2

@@ Abstract
We study Bayesian inverse problems with mixed noise, modeled as a combination of additive and multiplicative Gaussian components. While traditional inference methods often assume fixed or known noise characteristics, real-world applications, particularly in physics and chemistry, frequently involve noise with unknown and heterogeneous structure. Motivated by recent advances in flow-based generative modeling, we propose a novel inference framework based on conditional flow matching embedded within an Expectation-Maximization (EM) algorithm to jointly estimate posterior samplers and noise parameters. To enable high-dimensional inference and improve scalability, we use simulation-free ODE-based flow matching as the generative model in the E-step of the EM algorithm. We find a more Bayesian interpretation of EM convergence conditions, and verify some slightly weaker ones on the mixed noise model. Our numerical results illustrate the effectiveness of combining EM inference with flow matching for mixed-noise Bayesian inverse problems.
@@ 1 Introduction
where $F\colon\mathbb{R}^{d}\to\mathbb{R}^{n}$ is a known possibly nonlinear forward operator and $\eta_{\theta}$ represents random noise depending on parameters $\theta$. In many real-world settings, the noise is not known, but depends on auxiliary noise parameters $\theta$, which are in our case two scalars $(a,b)$. We focus on so-called mixed noise conditional on $X=x$, which is a sum of additive and multiplicative Gaussian noise, i.e.,
@@
As many physical phenomena (or approximations thereof) do not usually exhibit Gaussian errors, this often yields a more flexible modelling approach. It is present in optics [17, 16] and also biology [27]. Furthermore, the mixed noise is a natural setting to extend to beyond additive noise that is still tractable for mathematical analysis.
In [14], an iterative procedure was introduced to jointly update normalizing flows and noise parameters, formulated within the Expectation-Maximization (EM) framework consisting of an so called $E$-step (expectation step) and a $M$-step (maximization step) [11]. This approach can be interpreted as a conditional extension of the DeepGEM algorithm [12]. However, the only convergence guarantee provided was the non-decreasing behavior of the marginal log-likelihood across iterations. Moreover, the method relies on discrete normalizing flows [28], which have recently been surpassed in performance and scalability by ODE-based normalizing flows [9], particularly those trained in a simulation-free manner via flow matching [23, 25]. Flow matching has shown great success, in particular in image generation and regularization [26], and is a variant of score-based diffusion [29] which does not rely on a standard Gaussian latent. This enables applications such as DNA sequence design, see [31]. Further, we can also make the latent distribution dependent on the data distribution, which yields optimal transport inspired flow matching approaches [8, 34].
@@
  • formulate an EM algorithm where the E-step is realized via flow matching, a modern generative modeling and sampling technique based on neural ODEs;
@@
  • establish consistency guarantees: in the population limit of infinitely many measurements generated under a true noise parameter $\theta^{*}$, we investigate under which criteria the EM procedure converges, and give plausibility arguments for the mixed noise model.
@@
We denote by $X\in\mathbb{R}^{d}$ the latent random variable, by $Y\in\mathbb{R}^{n}$ the observation and $\Theta\subset\mathbb{R}^{d}$ a nonempty, compact and convex parameter space. We denote by $\|\cdot\|$ the Euclidean norm on $\mathbb{R}^{d}$. For a function $\mathcal{Q}\colon\Theta\times\Theta\to\mathbb{R}$, we denote by $\nabla_{1}$ the gradient with respect to the first input.
Assuming existence (cp. [32]), we denote by $p_{Y_{\theta}}$ the density of the marginal likelihood of $Y$ given $\theta$, by $p_{X|Y_{\theta}=y}$ the corresponding posterior density and by $p_{X,Y_{\theta}}$ the joint density. The corresponding distributions are denoted using capitalization, e.g. $P_{X|Y_{\theta}=y}$ for the posterior distribution. We denote by $\mathrm{pdf}(\mathbb{R}^{d})$ the set of all Lebesgue densities on $\mathbb{R}^{d}$, i.e.,
@@
whenever $\mu$ is absolutely continuous with respect to $\nu$ and the integral is finite and set to $\infty$ else. For measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$ with Lebesque densities $f,g\in\mathrm{pdf}(\mathbb{R}^{d})$ we may use abuse of notation and define
@@
Project Context
The project can be seen in the context of the PhD project of the first author, supervised by the last author. In particular, it falls into the category of how one can solve inverse problems using conditional generative models. The main theme is that one can learn joint maps, i.e., families of generative models parameterized by the observation of the inverse problem. This is a very common theme in machine learning and for instance done in [5, 3, 14]. This paper, in its theory also builds upon the classical Bayesian stability theory [20, 30, 32] which was subsequently used to obtain pointwise estimates for conditional generators [2] within this project. Therefore this paper revolves around the themes of the PhD project.
Outline
The manuscript at hand is organized as follows. We start giving the relevant background information on the general EM algorithm and flow matching in Section˜2.1 and Section˜2.2 respectively. Next, we recall the convergence analysis of the so-called population EM operator in Section˜3 and extend the theory in the follow up subsections Section˜3.3 and Section˜3.4. Then, we will verify the new framework for a class of mixed noise models in Section˜4. Finally, we illustrate the proposed method of mixed-noise learning with flow-matching in the numerical section for MNIST denoising in Section˜5.1 and for a reconstruction task with underlying partial differential equation (PDE) in Section˜5.2.
@@ 2.1 EM Algorithm
For the E-step, classical approaches include MCMC-based posterior sampling. However, such methods are computationally demanding, especially when repeated for each observation. Alternatively, we adopt a conditional generative modeling perspective, where a single shared model is trained to approximate the posterior distribution for all observations. This amortized setup enables efficient inference across the dataset and allows parameter sharing across different $y_{i}$. While [14] relied on conditional normalizing flows in the spirit of [12], we employ conditional flow matching as introduced in [1, 23, 24], which provides improved stability and scalability in high-dimensional settings (which can be seen by improved image generation results as opposed to discrete normalizing flows).
@@ 2.2 Flow Matching
We now detail the concrete realization of the E-step in our EM algorithm using conditional flow matching. The E-step aims to approximate the posterior distributions
@@
for each observation $y_{i}$, given current parameter estimates $\theta^{(k)}$. In [14], these $q_{i}$ were approximated using conditional normalizing flows, extending the DeepGEM framework [12] to the case of unknown noise parameters.
@@ 3 Convergence Analysis of the Population EM Operator
We start by recalling the convergence results for the population EM algorithm from [4]. For $\theta\in\Theta$ consider a parametric family of distributions $P_{Y_{\theta}}$ with density $p_{Y_{\theta}}$. We assume that the observed data are generated according to a distribution $P_{Y_{\theta^{*}}}$ with density $p_{Y_{\theta^{*}}}$ corresponding to a true but unknown parameter $\theta^{*}\in\Theta$. More precisely, we assume that $\theta^{*}$ is the unique maximizer of the population log-likelihood
$\theta^{*}=\operatorname*{arg\,max}_{\theta\in\Theta}\,\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}[\log p_{Y_{\theta}}(y)].$
@@
$\mathcal{Q}(\theta,\hat{\theta}):=\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}\,\mathbb{E}_{x\sim p_{X\mid Y_{\hat{\theta}}=y}}\left[\log p_{X,Y_{\theta}}(x,y)\right].$
@@ 3.1 Sufficient Assumptions
In this subsection, we are interested in deriving sufficient conditions for two scenarios. First we focus on conditions that yield local strong concavity and first order stability for the case of population EM iteration. However, as it will turn out, the bound in the Wasserstein case will be too rough to give convergence, therefore we retreat to showing the FOS condition with high probability.
@@
    (Smoothness and Integrability of Log-Likelihood) The map $\theta\mapsto\log p_{X,Y_{\theta}}$is two times continuously differentiable in $B_{\varepsilon}(\theta^{*})$ for all $(x,y)\in\mathbb{R}^{d\times n}$ and some $\varepsilon>0$ and there exist functions $G,H:\mathbb{R}^{d\times n}\to\mathbb{R}$ such that for all $(x,y)\in\mathbb{R}^{d\times n}$ and all $\theta\in B_{\varepsilon}(\theta^{*})$ it holds
@@
    $\int_{\mathbb{R}^{d\times n}}G(x,y)\,p_{X|Y_{\theta}=y}\,p_{Y_{\theta^{*}}}\text{ d}x\text{ d}y<C,\quad\int_{\mathbb{R}^{d\times n}}H(x,y)\,p_{X|Y_{\theta}=y}\,p_{Y_{\theta^{*}}}\text{ d}x\text{ d}y<C.$
@@
    (Local identifiability of $\theta^{\ast}$) The Fisher information matrix
    $\mathcal{I}(\theta^{\ast})\coloneqq-\mathbb{E}_{x,y}[\nabla^{2}_{\theta}\log p_{X,Y_{\theta^{\ast}}}]$
    satisfies $\mathcal{I}(\theta^{*})\geq\lambda_{\ast}I_{d}$, for some $\lambda_{\ast}>0$ and denote the smallest constant $\mu^{\ast}$ with $\mathcal{I}(\theta^{*})\leq\mu^{\ast}I_{d}$ with $\mu^{\ast}>0$.
@@
    The map $\theta\mapsto p_{X|Y_{\theta}=y}$ is locally $L(y)$- Lipschitz in the Wasserstein-1 distance, i.e. for all $(x,y)\in\mathbb{R}^{d\times n}$
@@
    $\left\|\nabla_{\theta}\log p_{X,Y_{\theta}}(x_{1},y)-\nabla_{\theta}\log p_{X,Y_{\theta}}(x_{2},y)\right\|\leq L_{g}(y,\theta)\|x_{1}-x_{2}\|\quad\forall y\in\mathbb{R}^{n},x_{1},x_{2}\in\mathbb{R}^{d}$
@@
    $\gamma=\gamma(\varepsilon)=\sup\limits_{\theta\in B_{\varepsilon}(\theta^{*})}\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}[L(y)L_{g}(\theta,y)]<\infty.$
@@
    $\gamma=\gamma(\varepsilon)=\sup\limits_{\theta\in B_{\varepsilon}(\theta^{*})}\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}[L(y)L_{g}({M}(\theta),y)]<\infty.$
  • (A5-SG)
    $Y$ is sub-Gaussian with constant $K$ [35], further there exist two continuous functions $h_{1},h_{2}$, such that the constants $L(Y)\leq h_{1}(\|Y\|)$ and $L_{g}(\theta,Y)\leq h_{2}(\|Y\|)$ for each $\theta\in B_{\varepsilon}(\theta^{*})$.
@@ 3.2 FOS Framework
Let (A1) and (A2) be fulfilled. Then there exist $\varepsilon>0$ and $\lambda=\lambda(\varepsilon,\lambda_{\ast})>0$ such that $\mathcal{Q}(\cdot,\theta^{\ast})$ is strong $\lambda$-concave on $B_{\varepsilon}(\theta^{*})$.
@@
Under regularity assumptions of (A1) we can swap the following differentiation and integration, and applying subsequently (A2), we get
$\nabla_{1}^{2}\mathcal{Q}(\theta^{\ast},\theta^{\ast})=\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}\mathbb{E}_{x\sim P_{X|Y_{\theta^{*}}=y}}\left[\nabla_{\theta}^{2}\log p_{X,Y_{\theta^{\ast}}}(x,y)\right]=-\mathcal{I}(\theta^{\ast})\leq-\lambda_{\ast}I_{d}.\\$ (5)
@@
for some $\lambda=\lambda(\varepsilon,\lambda_{\ast})>0$, which yields the strong $\lambda$-concavity $\mathcal{Q}(\cdot,\theta^{\ast})$. ∎
@@
$\nabla_{1}\mathcal{Q}(\theta,\hat{\theta})-\nabla_{1}\mathcal{Q}(\theta,\theta^{*})$ $=\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}\Big[\mathbb{E}_{x\sim P_{X|Y_{\hat{\theta}}=y}}\left[\nabla_{\theta}\log p_{X,Y_{\theta}}(x,y)\right]$
@@
$=\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}\Big[\mathbb{E}_{x\sim P_{X|Y_{\hat{\theta}}=y}}\left[\phi_{y,\theta}(x)\right]-\mathbb{E}_{x\sim P_{X|Y_{\theta^{*}}=y}}\left[\phi_{y,\theta}(x)\right]\Big].$
@@
Integrating over $y\sim p_{Y_{\theta^{*}}}$, we obtain
$\left\|\nabla_{1}\mathcal{Q}(\theta,\hat{\theta})-\nabla_{1}\mathcal{Q}(\theta,\theta^{*})\right\|\leq\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}[L(y)L_{g}(\theta,y)]\|\hat{\theta}-\theta^{\ast}\|.$
@@
$\left\|\nabla_{1}\mathcal{Q}(M(\hat{\theta}),\hat{\theta})-\nabla_{1}\mathcal{Q}(M(\hat{\theta}),\theta^{*})\right\|$ $\leq\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}[L(y)L_{g}(M(\hat{\theta}),y)]\|\hat{\theta}-\theta^{\ast}\|$
@@
Hence, we get slightly different assumptions: The score $\nabla_{\theta}\log p_{X,Y_{\theta}}$ now only needs to be element of $L^{\infty}(\mathbb{R}^{d})$ for all $y$, whereas the Lipschitz continuity bound is a bit harder to satisfy.
Note that in order to apply the convergence theorem, we must assume that $0\leq\gamma<\lambda$. There are two main reasons why we prefer the following gradient-based framework over the classical EM convergence theory that relies on the FOS condition. First, the FOS condition involves the population EM operator $M$, whose analytical form is often intractable and difficult to estimate numerically, especially in complex models. Second, Lemma 9 provides a slightly improved convergence guarantee: instead of requiring $\gamma<\lambda$, it suffices to assume $\gamma<c$, where $c$ is the strong convexity constant of the Q-function. This is a strictly weaker condition in non-degenerate cases. Therefore, we adopt the gradient-based framework, which is both more natural for our approximate M-step and yields more accessible convergence criteria.
As it will turn out, the integrability of $\gamma$ will not be true for our mixed noise model. However, luckily assumption (A5SG) will help us to show a similar statement with high probability. For this we look at the FOS condition in a probabilistic way instead, which means instead of bounding the expectation, we get a corresponding high probability statement. To rigorously formulate it, recall the definition of the $Q$-function and define $q$ via
$\mathcal{Q}(\theta,\hat{\theta})=\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}\,\mathbb{E}_{x\sim p_{X\mid Y_{\hat{\theta}}=y}}\left[\log p_{X,Y_{\theta}}(x,y)\right]=\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}[q(y,\hat{\theta},\theta)].$
Proposition 6.
Let (A1), (A3),(A4) and (A5SG) be fulfilled. For each $0<\delta\leq 1$, there exists a measurable subset $\Omega_{\delta}\subset\Omega$ (where $(\Omega,\mathcal{F},\mathbb{P})$ is the underlying probability space) and a constant $C_{\delta}>0$ such that
$\|\nabla_{\theta}q(Y(\omega),\theta^{*},\theta)-\nabla_{\theta}q(Y(\omega),\hat{\theta},\theta)\|\leq C_{\delta}\|\theta^{*}-\hat{\theta}\|,\qquad\forall\omega\in\Omega_{\delta},\theta\in B_{\varepsilon}(\theta^{*})$
holds with probability $\mathbb{P}(\Omega_{\delta})\geq 1-\delta$.
Proof.
As in the proof of Proposition 2, we can bound,
$\|\nabla_{\theta}q(y,\theta^{*},\theta)-\nabla_{\theta}q(y,\hat{\theta},\theta)\|$ $\leq L(y)L_{g}(\theta,y)\|\hat{\theta}-\theta^{\ast}\|.$
The sub-Gaussianity from (A5SG) yields a bound
$\mathbb{P}(\|Y\|\geq t)\leq 2\exp(-t^{2}/K).$
Via the continuity of the functions $h_{1}$ and $h_{2}$, we can then bound $L(Y)$ and $L_{g}(\theta,Y)$ with high probability as required in the statement. ∎